OpenStudy (anonymous):

-6x^2+4x-9=0 (What is the value of the discrimiant)

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]\[-6x^2+4x-9=0\] \[\Delta =b^2-4ac\]